OpenStudy (anonymous):

2x*f(x)-f(1-x)=2x^3+3x^2+4x-2 f function defined with reel numbers. f(1)+[dy/dx f(1)]=?

OpenStudy (anonymous):

hm.... how do you get the f(1-x) in terms of f(x)

OpenStudy (anonymous):

You can try x = 0, plug it in and solve for f(1)

OpenStudy (anonymous):

So 0 = -2 + f(1)

OpenStudy (anonymous):

f(1) = 2

OpenStudy (anonymous):

any suggestions.. what class is this for? calc 3?

OpenStudy (anonymous):

I could find f(1) , but can't find dy/dx f(1) , it's a derivative question.

OpenStudy (anonymous):

I know... hmm...

OpenStudy (anonymous):

Oh.. use implict deriavtion... and do the same thing by replacing f(0).

OpenStudy (anonymous):

derivated case of equation is , 2*f(x)+2x*(dy/dx f(x))+(dy/dx f(1-x))=6x^2+6x+4 , if we plug in 0 , we get f(0) and dy/dx f(0) , how can we get rid of these?

OpenStudy (anonymous):

If you plug in 0, it will cancel the dy/dx f(x) and left you with 2*f(0) + dydxf(1-x) = 4.. Now you try to figure out f(0) from the regular equation.

OpenStudy (anonymous):

cheers anyway...